OpenStudy (anonymous):

2. If x © y= (x+y) ^2/(x-y) ^2 then which of the following equals 1? A. 1© -1 B. 2©1 C. 2©2 D. -1©0 E.none of these

OpenStudy (anonymous):

Just plugging in the values of x and y, D.: (-1 + 0)^2 = 1 (-1 - 0)^2 = 1 1/1 = 1

OpenStudy (anonymous):

So the answer is D .Is it right

OpenStudy (anonymous):

I think yeah, but wait for a second opinion.